# web_appendix

### v4

@@ 4.1 Target parameter
$\delta_{2}:=$ $E\left(\frac{|D_{2}-D_{1}|}{E(|D_{2}-D_{1}||S=1)}\times\frac{Y_{2}(D_{2})-Y_{2}(D_{1})}{D_{2}-D_{1}}\middle|S=1\right)$
@@
$=$ $\frac{E\left(\text{sgn}(D_{2}-D_{1})(Y_{2}(D_{2})-Y_{2}(D_{1}))\right)}{E(|D_{2}-D_{1}|)}.$
$\delta_{2}$ is a weighted average of the slopes of switchers’ potential outcome functions from their period-one to their period-two treatments, where slopes receive a weight proportional to switchers’ absolute treatment change from period one to two. Accordingly, we refer to $\delta_{2}$ as the Weighted Average of Slopes (WAS). $\delta_{2}=\delta_{1}$ if and only if
$\text{cov}\left(\frac{Y_{2}(D_{2})-Y_{2}(D_{1})}{D_{2}-D_{1}},|D_{2}-D_{1}|\middle|S=1\right)=0:$ (3)
the WAS and AS are equal if and only if switchers’ slopes are uncorrelated with $|D_{2}-D_{1}|$.
Economically, the AS and WAS serve different purposes. As discussed above, under shape restrictions on the potential outcome function, the AS can be used to identify or bound the effect of other treatment changes than the actual change switchers experienced from period one to two. The WAS cannot serve that purpose, but under some assumptions, it may be used to conduct a cost-benefit analysis of the treatment changes that took place from period one to two. To simplify the discussion, let us assume in the remainder of this paragraph that $D_{2}\geq D_{1}$. Assume also that the outcome is a measure of output, such as agricultural yields or wages, expressed in monetary units. Finally, assume that the treatment is costly, with a cost linear in dose, uniform across units, and known to the analyst: the cost of giving $d$ units of treatment to a unit at period $t$ is $c_{t}\times d$ for some known $(c_{t})_{t\in\{1,2\}}$. Then, $D_{2}$ is beneficial relative to $D_{1}$ if and only if $E(Y_{2}(D_{2})-c_{2}D_{2})>E(Y_{2}(D_{1})-c_{2}D_{1})$ or, equivalently,
$\delta_{2}>c_{2}:$
comparing $\delta_{2}$ to $c_{2}$ is sufficient to evaluate if changing the treatment from $D_{1}$ to $D_{2}$ was beneficial.
4.2 Identification
Let $S_{+}=1\{D_{2}-D_{1}>0\}$, $S_{-}=1\{D_{2}-D_{1}<0\}$ and
$\delta_{2+}$ $:=\frac{E\left(Y_{2}(D_{2})-Y_{2}(D_{1})|S_{+}=1\right)}{E(D_{2}-D_{1}|S_{+}=1)},$
$\delta_{2-}$ $:=\frac{E\left(Y_{2}(D_{1})-Y_{2}(D_{2})|S_{-}=1\right)}{E(D_{1}-D_{2}|S_{-}=1)}.$
Hereafter, units with $S_{+}=1$ are referred to as “switchers up”, while units with $S_{-}=1$ are referred to as “switchers down”. Thus, $\delta_{2+}$ is the WAS of switchers up, and $\delta_{2-}$ is the WAS of switchers down. One has
$\delta_{2}=$ $\frac{P(S_{+}=1|S=1)E(D_{2}-D_{1}|S_{+}=1)}{E(|D_{2}-D_{1}||S=1)}\delta_{2+}$
$+$ $\frac{P(S_{-}=1|S=1)E(D_{1}-D_{2}|S_{-}=1)}{E(|D_{2}-D_{1}||S=1)}\delta_{2-}.$ (4)
To identify $\delta_{2+}$ (resp. $\delta_{2-}$) we use DID estimands comparing switchers up (resp. switchers down) to stayers with the same period-one treatment. This requires that there be no value of $D_{1}$ such that some switchers up (resp. switchers down) have that baseline treatment while there is no stayer with the same baseline treatment, as stated in Point 1 (resp. 2) of Assumption 7 below.
@@ 4.2 Identification
(Support conditions for WAS identification)
  1. 1.
    $0<P(S_{+}=1)$, and $0<P(S_{+}=1|D_{1})$ implies that $0<P(S=0|D_{1})$.
  2. 2.
    $0<P(S_{-}=1)$, and $0<P(S_{-}=1|D_{1})$ implies that $0<P(S=0|D_{1})$.
@@
  1. 1.
    If Assumptions 1-2 and Point 1 of Assumption 7 hold,
    $\delta_{2+}$ $=\frac{E\left(\Delta Y-E(\Delta Y|D_{1},S=0)\middle|S_{+}=1\right)}{E(\Delta D|S_{+}=1)}$ (5)
    $=\frac{E\left(\Delta Y|S_{+}=1\right)-E\left(\Delta Y\frac{P(S_{+}=1|D_{1})}{P(S=0|D_{1})}\frac{P(S=0)}{P(S_{+}=1)}\middle|S=0\right)}{E(\Delta D|S_{+}=1)}.$ (6)
  2. 2.
    If Assumptions 1-2 and Point 2 of Assumption 7 hold,
    $\delta_{2-}$ $=\frac{E\left(\Delta Y-E(\Delta Y|D_{1},S=0)\middle|S_{-}=1\right)}{E(\Delta D|S_{-}=1)}$ (7)
    $=\frac{E\left(\Delta Y|S_{-}=1\right)-E\left(\Delta Y\frac{P(S_{-}=1|D_{1})}{P(S=0|D_{1})}\frac{P(S=0)}{P(S_{-}=1)}\middle|S=0\right)}{E(\Delta D|S_{-}=1)}.$ (8)
  3. 3.
    If Assumptions 1-2 and Assumption 7 hold,
    $\delta_{2}=$ $\frac{E\left[\text{sgn}(\Delta D)\left(\Delta Y-E(\Delta Y|D_{1},S=0)\right)\right]}{E[|\Delta D|]}$ (9)
    $=$ $\frac{E\left[\text{sgn}(\Delta D)\Delta Y\right]-E\left[\Delta Y\frac{P(S_{+}=1|D_{1})-P(S_{-}=1|D_{1})}{P(S=0|D_{1})}P(S=0)\middle|S=0\right]}{E[|\Delta D|]}.$ (10)
Point 1 of Theorem 4 shows that $\delta_{2+}$, the WAS of switchers-up, is identified by two estimands, a regression-based and a propensity-score-based estimand. Point 2 of Theorem 4 shows that $\delta_{2-}$, the WAS of switchers down, is identified by two estimands similar to those identifying $\delta_{2+}$, replacing $S_{+}$ by $S_{-}$. Finally, if the conditions in Point 1 and 2 of Theorem 4 jointly hold, it directly follows from (4.2) that $\delta_{2}$, the WAS of all switchers, is identified by a weighted average of the estimands in Equations (5) and (7), and by a weighted average of the estimands in Equations (6) and (8). Those weighted averages simplify into the expressions given in Point 3 of Theorem 4. Point 3 of Theorem 4 also implies that $\delta_{2}$ is identified by the following doubly-robust estimand:
$\frac{E\left[\left(S_{+}-S_{-}-\frac{P(S_{+}=1|D_{1})-P(S_{-}=1|D_{1})}{P(S=0|D_{1})}(1-S)\right)(\Delta Y-E(\Delta Y|D_{1},S=0))\right]}{E[\left|\Delta D\right|]}.$ (11)
4.3 Estimation and inference
The regression-based estimands identifying $\delta_{2+}$ and $\delta_{2-}$ can be estimated following almost the same steps as in Section 3.3. Specifically, let
$\widehat{\delta}^{r}_{2+}:=$ $\frac{\frac{1}{n_{+}}\sum_{i:S_{i+}=1}\left(\Delta Y_{i}-\widehat{E}(\Delta Y|D_{1}=D_{i,1},S=0)\right)}{\frac{1}{n_{+}}\sum_{i:S_{i+}=1}\Delta D_{i}}$
$\widehat{\delta}^{r}_{2-}:=$ $\frac{\frac{1}{n_{-}}\sum_{i:S_{i-}=1}\left(\Delta Y_{i}-\widehat{E}(\Delta Y|D_{1}=D_{i,1},S=0)\right)}{\frac{1}{n_{-}}\sum_{i:S_{i-}=1}\Delta D_{i}},$
where $n_{+}=\#\{i:S_{i+}=1\}$ and $n_{-}=\#\{i:S_{i-}=1\}$, and where $\widehat{E}(\Delta Y|D_{1},S=0)$ is the series estimator of $E(\Delta Y|D_{1},S=0)$ defined in Section 3.3 of the paper. Then, let
$\widehat{w}_{+}=\frac{\frac{n_{+}}{n}\times\frac{1}{n_{+}}\sum_{i:S_{i+}=1}\Delta D_{i}}{\frac{n_{+}}{n}\times\frac{1}{n_{+}}\sum_{i:S_{i+}=1}\Delta D_{i}-\frac{n_{-}}{n}\times\frac{1}{n_{-}}\sum_{i:S_{i-}=1}\Delta D_{i}},$
and let
$\widehat{\delta}^{r}_{2}=$ $\widehat{w}_{+}\widehat{\delta}^{r}_{2+}+(1-\widehat{w}_{+})\widehat{\delta}^{r}_{2-}$
be the corresponding estimator of $\delta_{2}$.
We now propose estimators of the propensity-score-based estimands identifying $\delta_{2+}$ and $\delta_{2-}$ in Equations (6) and (8). Let $\widehat{P}(S_{+}=1)=n_{+}/n$ (resp. $\widehat{P}(S_{-}=1)=n_{-}/n$, $\widehat{P}(S=0)=(n-n_{s})/n$) be an estimator of $P(S_{+}=1)$ (resp. $P(S_{-}=1)$, $P(S=0)$). Let $\widehat{P}(S_{+}=1|D_{1})$ (resp. $\widehat{P}(S_{-}=1|D_{1})$, $\widehat{P}(S=0|D_{1})$) be a non-parametric estimator of $P(S_{+}=1|D_{1})$ (resp. $P(S_{-}=1|D_{1})$, $P(S=0|D_{1})$) using a series logistic regression of $S_{i+}$ (resp. $S_{i-}$, $1-S_{i}$) on polynomials in $D_{1}$ $(p_{k,K_{n}}(D_{1}))_{1\leq k\leq K_{n}}$. We make the following technical assumption.
Assumption 8
(Conditions for asymptotic normality of propensity-score WAS estimator)
  1. 1.
    $D_{1}$ is continuously distributed on a compact interval $I$, with $\inf_{d\in I}f_{D_{1}}(d)>0$.
  2. 2.
    $E[\Delta Y^{2}]<\infty$ and $d\mapsto E[\Delta Y^{2}|D_{1}=d]$ is bounded on $I$
  3. 3.
    $10<E[S_{+}]<$ 1, $10<E[S_{-}]<$ 1, $E[S]>0$ and $\sup_{d\in I}E[S|D_{1}=d]<1$.
  4. 4.
    The functions $d\mapsto E[\Delta Y(1-S)|D_{1}=d]$, $d\mapsto E[S|D_{1}=d]$, $d\mapsto E[S_{+}|D_{1}=d]$ and $d\mapsto E[S_{-}|D_{1}=d]$ are four times continuously differentiable.
  5. 5.
    The polynomials $d\mapsto p_{k,K_{n}}(d)$, $k\leq 1\leq K_{n}$ are orthonormal on $I$ and $K_{n}=Cn^{\nu}$ where $61/10<\nu<1/$ 6.
Let
$\widehat{\delta}^{ps}_{2+}:=$ $\frac{\frac{1}{n_{+}}\sum_{i:S_{i+}=1}\Delta Y_{i}-\frac{1}{n-n_{s}}\sum_{i:S_{i}=0}\Delta Y_{i}\frac{\widehat{P}(S_{+}=1|D_{1}=D_{i1})}{\widehat{P}(S=0|D_{1}=D_{i1})}\frac{\widehat{P}(S=0)}{\widehat{P}(S_{+}=1)}}{\frac{1}{n_{+}}\sum_{i:S_{i+}=1}\Delta D_{i}}$
$\widehat{\delta}^{ps}_{2-}:=$ $\frac{\frac{1}{n_{-}}\sum_{i:S_{i-}=1}\Delta Y_{i}-\frac{1}{n-n_{s}}\sum_{i:S_{i}=0}\Delta Y_{i}\frac{\widehat{P}(S_{-}=1|D_{1}=D_{i1})}{\widehat{P}(S=0|D_{1}=D_{i1})}\frac{\widehat{P}(S=0)}{\widehat{P}(S_{-}=1)}}{\frac{1}{n_{-}}\sum_{i:S_{i-}=1}\Delta D_{i}},$
and let
$\widehat{\delta}^{ps}_{2}=$ $\widehat{w}_{+}\widehat{\delta}^{ps}_{2+}+(1-\widehat{w}_{+})\widehat{\delta}^{ps}_{2-}$
be the corresponding estimator of $\delta_{2}$. Let
$\psi_{2+}$ $:=\frac{1}{E(\Delta DS_{+})}\left\{\left(S_{+}-E(S_{+}|D_{1})\frac{(1-S)}{E(1-S|D_{1})}\right)(\Delta Y-E(\Delta Y|D_{1},S=0))-\delta_{2+}\Delta DS_{+}\right\}$
$\psi_{2-}$ $:=\frac{1}{E(\Delta DS_{-})}\left\{\left(S_{-}-E(S_{-}|D_{1})\frac{(1-S)}{E(1-S|D_{1})}\right)(\Delta Y-E(\Delta Y|D_{1},S=0))-\delta_{2-}\Delta DS_{-}\right\}$
$\psi_{2}$ $:=\frac{1}{E(\left|\Delta D\right|)}\left\{\left(S_{+}-S_{-}-E(S_{+}-S_{-}|D_{1})\frac{(1-S)}{E(1-S|D_{1})}\right)\right.$
×$\qquad\times(\Delta Y-E(\Delta Y|D_{1},S=0))-\delta_{2}\left|\Delta D\right|\bigg{\}}.$
Theorem 5
  1. 1.
    If Assumptions 1-3 and 6 hold,
    $\sqrt{n}\left((\widehat{\delta}^{r}_{2+},\widehat{\delta}^{r}_{2-})^{\prime}-(\delta_{2+},\delta_{2-})^{\prime}\right)\stackrel{{\scriptstyle d}}{{\longrightarrow}}\mathcal{N}(0,V((\psi_{2+},\psi_{2-})^{\prime})).$
    and
@@ 4.3 Estimation and inference
    $\sqrt{n}\left(\widehat{\delta}^{r}_{2}-\delta_{2}\right)\stackrel{{\scriptstyle d}}{{\longrightarrow}}\mathcal{N}(0,V(\psi_{2})).$
  2. 2.
    If Assumptions 1-3 and 8 hold,
    $\sqrt{n}\left((\widehat{\delta}^{ps}_{2+},\widehat{\delta}^{ps}_{2-})^{\prime}-(\delta_{2+},\delta_{2-})^{\prime}\right)\stackrel{{\scriptstyle d}}{{\longrightarrow}}\mathcal{N}(0,V((\psi_{2+},\psi_{2-})^{\prime})).$
    and
    $\sqrt{n}\left(\widehat{\delta}^{ps}_{2}-\delta_{2}\right)\stackrel{{\scriptstyle d}}{{\longrightarrow}}\mathcal{N}(0,V(\psi_{2})).$
Based on (11), we can also estimate $\delta_{2}$ using the following doubly-robust estimator:
$\hat{\delta}^{dr}_{2}=\frac{\sum_{i}\left(S_{i+}-S_{i-}-\frac{\hat{P}(S_{+}=1|D_{1}=D_{1i})-P(S_{i-}=1|D_{1}=D_{1i})}{P(S_{i}=0|D_{1}=D_{1i})}(1-S_{i})\right)(\Delta Y_{i}-\hat{E}(\Delta Y_{i}|D_{1}=D_{1i},S_{i}=0))}{\sum_{i}\left|\Delta D_{i}\right|}.$
This estimator has an important advantage. While our regression-based (resp. propensity-score-based) estimator is nominally non-parametric, in practice it requires choosing a polynomial order to estimate $E(\Delta Y|D_{1},S=0)$ (resp. $P(S_{+}=1|D_{1})$ and $P(S_{-}=1|D_{1})$), and the rate conditions on $K_{n}$ in Assumptions 6 (resp. 8) do not give specific guidance on the choice of this tuning parameter. With the doubly-robust estimator above, one can choose this tuning parameter in a data-driven manner, using cross-validation (CV). Results in Section 4 of Andrews (1991) imply that a series estimator of a nonparametric regression model with a polynomial order chosen by CV is optimal: the ratio of its mean-squared error and that of an oracle estimator using the best polynomial order given the sample size converges to one. Then, as $D_{1}$ is a scalar variable, series estimators of $E(\Delta Y|D_{1},S=0)$, $P(S_{+}=1|D_{1})$, and $P(S_{-}=1|D_{1})$ with CV-chosen polynomial orders converge at a rate strictly faster than $n^{1/4}$, as long as one assumes that those nuisance functions are twice differentiable. Then, we conjecture that one can show, following arguments similar to those in Farrell (2015), that $\hat{\delta}^{dr}_{2}$ with CV-chosen polynomial orders in the estimation of the nuisance functions is $\sqrt{n}-$consistent, with asymptotic variance $V(\psi_{2})$.
@@
If Assumption 1 holds, $(Y_{2}(D_{2})-Y_{2}(D_{1}))/(D_{2}-D_{1})=\delta$ for some real number $\delta$, $V(\Delta Y(D_{1})|D_{1},D_{2})=\sigma^{2}$ for some real number $\sigma^{2}>0$, $D_{2}\geq D_{1}$, and $\Delta D\perp\!\!\!\perp D_{1}$,
@@
Of course, the constant treatment effect and the homoscedasticity assumptions underlying Proposition 1 are strong, but one often has to make strong assumptions to be able to rank estimators’ variances. The question then is whether this ranking still holds in real-life applications, where those assumptions are unlikely to hold. Put differently, all models are wrong but some are useful, and the question is whether Proposition 1 is useful. In our empirical application, we find that the variance of $\hat{\delta}_{1}$ is indeed much larger than that of $\hat{\delta}^{dr}_{2}$, as predicted by Proposition 1.
5 Instrumental-variable estimation
There are instances where the parallel-trends condition in Assumption 1 is implausible, but one has at hand an instrument satisfying a similar parallel-trends condition. For instance, one may be interested in estimating the price-elasticity of a good’s consumption, but prices respond to supply and demand shocks, and therefore do not satisfy Assumption 1. On the other hand, taxes may not respond to supply and demand shocks and may satisfy a parallel-trends assumption.
5.1 Notation and assumptions
Let $(Z_{1},Z_{2})$ denote the instrument’s values at period one and two and $\mathcal{Z}_{t}$ be the support of $Z_{t}$. For any $z\in\mathcal{Z}_{1}\cup\mathcal{Z}_{2}$, let $D_{1}(z)$ and $D_{2}(z)$ respectively denote the unit’s potential treatments at periods 1 and 2 with instrument $z$. Let $SC=1\{D_{2}(Z_{2})\neq D_{2}(Z_{1}),Z_{2}\neq Z_{1}\}$ be an indicator equal to 1 for switchers-compliers, namely units whose instrument changes from period one to two and whose treatment is affected by that change in the instrument.
We replace Assumption 1 by the following assumption. [^5] Note that with our notation where potential outcomes do not depend on $z$, we also implicitly impose the usual exclusion restriction.

### v5

@@ 4.1 Target parameter
$\delta_{2}:=$ $E\left(\frac{|D_{2}-D_{1}|}{E(|D_{2}-D_{1}||S=1)}\times\frac{Y_{2}(D_{2})-Y_{2}(D_{1})}{D_{2}-D_{1}}\middle|S=1\right).$
$\delta_{2}$ is a weighted average of the slopes of switchers’ potential outcome functions from their period-one to their period-two treatments, where slopes receive a weight proportional to switchers’ absolute treatment change from period one to two. Accordingly, we refer to $\delta_{2}$ as the Weighted Average of Slopes (WAS). The WAS and AS are equal if and only if switchers’ slopes are uncorrelated with $|D_{2}-D_{1}|$.
The AS and WAS serve different purposes. Unlike the AS, the WAS may not be used to identify or bound the effect of other treatment changes than the actual changes that took place, even under shape restrictions on the potential outcome function. Instead, the WAS may be used to conduct a cost-benefit analysis of the treatment changes that took place from period one to two. To simplify the discussion, let us assume in the remainder of this paragraph that $D_{2}\geq D_{1}$. Assume also that the outcome is a measure of output, such as agricultural yields or wages, expressed in monetary units. Finally, assume that the treatment is costly, with a cost linear in dose, uniform across units, and known to the analyst: the cost of giving $d$ units of treatment to a unit at period $t$ is $c_{t}\times d$ for some known $(c_{t})_{t\in\{1,2\}}$. Then, $D_{2}$ is beneficial relative to $D_{1}$ if and only if $E(Y_{2}(D_{2})-c_{2}D_{2})>E(Y_{2}(D_{1})-c_{2}D_{1})$, namely if and only if $\delta_{2}>c_{2}:$ comparing $\delta_{2}$ to $c_{2}$ is sufficient to evaluate if changing the treatment from $D_{1}$ to $D_{2}$ was beneficial.
4.2 Identification
Let $S_{+}=1\{D_{2}-D_{1}>0\}$ and $S_{-}=1\{D_{2}-D_{1}<0\}$ respectively be indicators for “switchers up” and “switchers down” and define the function:
$\gamma_{0}(s,d):=\frac{E(S_{+}|D_{1}=d)-E(S_{-}|D_{1}=d)}{E(1-S|D_{1}=d)}(1-s).$
Theorem 4
If Assumptions 1, 2, and 4 hold, then for any functions $(s,d)\mapsto\gamma(s,d)$ and $d\mapsto\mu(d)$, respectively defined on $\{0,1\}\times\mathcal{D}_{1}$ and $\mathcal{D}_{1}$ and such that $\gamma$ is bounded and $E[|\mu(D_{1})|]<\infty$:
$\delta_{2}=$ $\frac{E\left[\left(S_{+}-S_{-}-\gamma(S,D_{1})\right)(\Delta Y-\mu_{0}(D_{1}))\right]}{E[\left|\Delta D\right|]}$
$=$ $\frac{E\left[\left(S_{+}-S_{-}-\gamma_{0}(S,D_{1})\right)(\Delta Y-\mu(D_{1}))\right]}{E[\left|\Delta D\right|]}.$
Here are the steps leading to Theorem 4. One has
$\delta_{2}=$ $E\left(\frac{|D_{2}-D_{1}|}{E(|D_{2}-D_{1}||S=1)}\times\frac{Y_{2}(D_{2})-Y_{2}(D_{1})}{D_{2}-D_{1}}\middle|S=1\right)$
@@ 4.2 Identification
$\qquad$== $\frac{E\left((S_{+}-S_{-})(Y_{2}(D_{2})-Y_{2}(D_{1}))\right)}{E(|\Delta D|)}$
$=$ $\frac{E\left((S_{+}-S_{-})(\Delta Y-\Delta Y(D_{1}))\right)}{E(|\Delta D|)}$
$=$ $\frac{E\left((S_{+}-S_{-})(E(\Delta Y|D_{1},D_{2})-E(\Delta Y(D_{1})|D_{1},D_{2}))\right)}{E(|\Delta D|)}$
$=$ $\frac{E\left((S_{+}-S_{-})(E(\Delta Y|D_{1},D_{2})-E(\Delta Y|D_{1},S=0))\right)}{E(|\Delta D|)}$
$=$ $\frac{E\left[\left(S_{+}-S_{-}\right)(\Delta Y-E(\Delta Y|D_{1},S=0))\right]}{E[\left|\Delta D\right|]}$
$=$ $\frac{E\left[\left(S_{+}-S_{-}\right)(\Delta Y-\mu_{0}(D_{1}))\right]}{E[\left|\Delta D\right|]}.$
The fourth equality follows from adding and subtracting $Y_{1}(D_{1})$. The fifth equality follows from the law of iterated expectations. The sixth follows from Assumption 1 and the fact that $\Delta Y=\Delta Y(D_{1})$ for stayers. The seventh follows from the law of iterated expectations, and the last follows from the definition of $\mu_{0}$. The numerator of the estimand identifying $\delta_{2}$ in the previous display is a conditional DID comparing the outcome evolutions of switchers and stayers with the same $D_{1}$. Following similar steps, one can show that $\delta_{2}$ is also identified by
$\frac{E\left[\left(S_{+}-S_{-}\right)\Delta Y\right]-E\left[\frac{E(S_{+}|D_{1}=d)-E(S_{-}|D_{1}=d)}{E(1-S|D_{1}=d)}(1-S)\Delta Y\right]}{E[\left|\Delta D\right|]}=\frac{E\left[\left(S_{+}-S_{-}-\gamma_{0}(S,D_{1})\right)\Delta Y\right]}{E[\left|\Delta D\right|]},$
an estimand whose numerator is an unconditional DID with propensity score reweighting. Theorem 4 implies that an estimand combining those two estimands,
$\frac{E\left[\left(S_{+}-S_{-}-\gamma_{0}(S,D_{1})\right)(\Delta Y-\mu_{0}(D_{1}))\right]}{E[\left|\Delta D\right|]},$
also identifies $\delta_{2}$, and is doubly-robust in the functions $\mu_{0}(D_{1})$ and $\gamma_{0}(S,D_{1})$.
4.3 Estimation and inference
Let $h_{0}(d):=E[S_{+}-S_{-}|D_{1}=d]$. As before, split the sample into two subsamples $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$, and define:
$\hat{\delta}_{2,\mathsf{DR}}^{(1)}:=\frac{1}{\Delta_{S}^{(1)}}\sum_{i\in\mathcal{I}_{1}}\left(S_{i+}-S_{i-}-\frac{\hat{h}^{(2)}(D_{i,1})}{\hat{p}^{(2)}(D_{i,1})}(1-S_{i})\right)(\Delta Y_{i}-\hat{\mu}^{(2)}(D_{i,1})).$
where $\Delta_{S}^{(1)}:=\sum_{i\in\mathcal{I}_{1}}\left|\Delta D_{i}\right|$ and we use the convention that $\hat{\delta}_{2,\mathsf{DR}}^{(1)}=0$ if $\Delta_{S}^{(1)}=0$. The estimator $\hat{\delta}_{2,\mathsf{DR}}^{(2)}$ is constructed analogously, permuting the roles of $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$. Finally, let
$\hat{\delta}_{2,\mathsf{DR}}:=\frac{\Delta_{S}^{(1)}}{\Delta_{S}^{(1)}+\Delta_{S}^{(2)}}\hat{\delta}_{2,\mathsf{DR}}^{(1)}+\frac{\Delta_{S}^{(2)}}{\Delta_{S}^{(1)}+\Delta_{S}^{(2)}}\hat{\delta}_{2,\mathsf{DR}}^{(2)}.$
We impose the following rate conditions.
@@ 4.3 Estimation and inference
For $k=1,2$,
  1. (a)
    $\left\|\hat{h}-h_{0}\right\|_{2,\mathcal{I}_{k}}=o_{P}(1)$, $\left\|\hat{\mu}-\mu_{0}\right\|_{2,\mathcal{I}_{k}}=o_{P}(1)$, $\left\|\hat{p}-p_{0}\right\|_{k,\infty}=o_{P}(1)$.
  2. (b)
    $\sqrt{n}\left\|\hat{h}-h_{0}\right\|_{2,\mathcal{I}_{k}}\times\left\|\hat{\mu}-\mu_{0}\right\|_{2,\mathcal{I}_{k}}=o_{P}(1)$, $\sqrt{n}\left\|\hat{p}-p_{0}\right\|_{2,\mathcal{I}_{k}}\times\left\|\hat{\mu}-\mu_{0}\right\|_{2,\mathcal{I}_{k}}=o_{P}(1)$.
@@
If Assumptions 1-4, 6, and 8 hold, $\sqrt{n}\left(\widehat{\delta}_{2,\mathsf{DR}}-\delta_{2}\right)\stackrel{{\scriptstyle d}}{{\longrightarrow}}\mathcal{N}(0,V(\psi_{2}))$, where
$\psi_{2}$ $:=\frac{1}{E(\left|\Delta D\right|)}\left\{\left(S_{+}-S_{-}-E(S_{+}-S_{-}|D_{1})\frac{(1-S)}{E(1-S|D_{1})}\right)\right.(\Delta Y-E(\Delta Y|D_{1},S=0))-\delta_{2}\left|\Delta D\right|\bigg{\}}.$
@@
If Assumption 1 holds, $(Y_{2}(D_{2})-Y_{2}(D_{1}))/(D_{2}-D_{1})=\delta$, $V(\Delta Y(D_{1})|D_{1},D_{2})=\sigma^{2}$ for some real number $\sigma^{2}>0$, $D_{2}\geq D_{1}$, and $\Delta D\perp\!\!\!\perp D_{1}$, then
@@
The constant effect and homoscedasticity assumptions underlying Proposition 1 are admittedly strong, but ranking estimators’ variances typically requires such conditions. The question then is whether the ranking in Proposition 1 still holds in real-life applications, where those assumptions are likely to fail. In our empirical application, we find that the variance of $\widehat{\delta}_{1,\mathsf{DR}}$ is indeed much larger than that of $\widehat{\delta}_{2,\mathsf{DR}}$.
5 Extensions
We now briefly discuss several extensions, on which we provide more details in the web appendix.
5.1 Instrumental-variable estimators
There are instances where Assumption 1 is implausible, but one has an instrument that may satisfy a similar parallel-trends condition. This section sketches the extension of our estimators to the IV case. For further details, see Section LABEL:appendixsec:IV of the web appendix.
Let $Z_{t}$ denote the instrumental variable at period $t\in\{1,2\}$ and $D_{t}(z)$ denote the unit’s potential treatment at $t$ with instrument value $z$. Let $SC=1\{D_{2}(Z_{2})\neq D_{2}(Z_{1})\}$ be an indicator equal to 1 for switchers-compliers, namely units whose instrument changes from period one to two and whose treatment is affected by that change in the instrument.
We replace Assumption 1 by the following assumption. [^6] Our notation, where potential outcomes do not depend on $z$, implicitly imposes the usual exclusion restriction.